OpenStudy (anonymous):

Suppose that a father is 28.00 y older than his daughter. He wants to travel outward from Earth for 4.700 y and then back for another 4.700 y (both intervals as he measures them) such that he is then 28.00 y younger than his daughter. What constant speed parameter β (relative to Earth) is required? (Give your answer to at least four significant figures.)

OpenStudy (shamim):

Formula is t=t0/(root (1-v^2/c^2))

OpenStudy (shamim):

v=?

OpenStudy (shamim):

c=light velocity=3*10^8m/s

OpenStudy (shamim):

t0=4.7+4.7=?y

OpenStudy (shamim):

t=28+28+4.7+4.7=?y